OpenStudy (anonymous):

reduce this radical 8/ radical 5

OpenStudy (anonymous):

radical 8 x radical 5 = radical 40 / radical 5 x radical 5 = radical 40 / 5 = radcal ( 4 x 10 ) / 5 = 2 radical 10 / 5

OpenStudy (anonymous):

\[\frac{ 8 }{\sqrt{5} }\]

OpenStudy (anonymous):

this is what it looks like

OpenStudy (anonymous):

i thought the answer would be \[\frac{ 8\sqrt{5} }{ 5 }\]

OpenStudy (anonymous):

(8 * √5) / (√5 * √5 ) = 8√5 / 5 so you are right.